A company uses 81 circuit boards a day in a manufacturing process. The person who orders the boards follows this rule: Order when the amount on hand drops to 630 boards. Orders are delivered approximately 7 days after being placed. The delivery time is normal with a mean of 7 days and a standard deviation of .90 days. What is the probability that the supply of circuit boards will be exhausted before the order is received if boards are reordered when the amount on hand drops to 630 boards?
Answer #1

Reorder point = demand during lead time + safety stock

= lead time x demand per day + z x SDLT x demand per day

630 = 7x81 + z x 0.9x81

z = 63 / 72.9 = 0.8641

which corresponds to  service level of 0.806

Probability of stock out = 1-0.806 =0.194

which means that there is 19.4% chance that the boards will be exhausted.